OpenStudy (anonymous):

solve the equation 10/n^2-1+2n-5/n-1=2n+5/n+1

OpenStudy (campbell_st):

I'll assume the question is \[\frac{10}{n^2 -1} + \frac{2n - 5}{n -1} = \frac{2n + 5}{n + 1}\] eliminate the denominator by multiplying every term by (n-1)(n+1) \[10 + (2n -5)(n+1) = (2n +5)(n -1)\] simplify the expression \[10 + 2n^2 -3n - 5 = 2n^2 + 3n - 5\] collect like terms gives \[10 - 6n = 0\] so 6n = 10 n = 5/3

OpenStudy (anonymous):

can you show how you eliminated the denominators?

OpenStudy (campbell_st):

ok... N^2 - 1 = (n-1)(n+1) 1st term \[\frac{10}{(n -1)(n+1)} \times (n -1)(n+1) = 10\] 2nd term \[\frac{(2n - 5)}{n -1} \times (n-1)(n+1) = (2n -5)(n +1)\] 3rd term \[\frac{(2n +5}{(n +1)} \times (n -1)(n +1) = (2n + 5)(n -1)\]

OpenStudy (anonymous):

oooh thank you very much it now makes more sense

OpenStudy (anonymous):

okay can you help me with this one 1/b+2+1/b+2=3/b+1

OpenStudy (campbell_st):

is this the problem..? \[\frac{1}{b +2}+ \frac{1}{b +2} = \frac{3}{b +1}\]

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

ok the left hand side can be simplifed \[\frac{2}{b +2} = \frac{3}{b +1}\] does that make sense

OpenStudy (anonymous):

yea i got that part, so do you eliminate the denominator by multiplying by b+2

OpenStudy (campbell_st):

just cross multiply to eliminate the denominators \[2(b +1) = 3(b +2)\]

OpenStudy (campbell_st):

really you multiply by (b+2)(b+1) but its the same as cross multiplying

OpenStudy (anonymous):

oh okay thanks again